# Efficient Reinforcement Learning for Optimal Control with Natural Images

### v1

@@ 3.6 Benchmark Performance with Alternative Sampling Strategies
Now we consider sampling strategies where only some of the states (image patches) of the benchmark are used to train the network. One possibility would be to focus training on initial states that have a suboptimal greedy policy. However, under an optimal policy states outside this set are visited. A better approach makes use of the findings in the previous section. Sampling from states with at least one non-negative coordinate for $c_{k}-t_{k}$ in the benchmark corresponds to training on the set of states where the tracker is not both behind and to the left of the target. Many of these states are independent of the set of states where the tracker is behind and to the left of the target: in this case the tracker never reaches the target. The result is a partition of the benchmark states. When states partition in this way, it is possible to focus training on the relevant partition to achieve optimal performance, as demonstrated in Fig 10.
@@
Figure 10: Expected total cost of optimal and greedy policies for each initial state with a suboptimal greedy policy when only 78% of the training set is used (see Text for details). Initial states are indexed from 1 to 672. Fitted VI remains optimal (Inset). The benchmark horizon is 400 time periods, and $p=0.4$.
@@
In Fig 10, the benchmark is used with a ×$\times 4$ sparse code leading to 1323 states (a small number of states is used here as it is easier to visualize graphically). Of these, 1031 states have at least one non-negative coordinate (i.e., 78% of the total number of states). These states are used as the training set. Fitted value iteration is then applied to find optimal policies for the 672 initial states with suboptimal greedy policies. Fig 10 shows these policies remain optimal policies under the proposed sampling strategy. However, decreasing the training set below 78% of the total number of states rapidly degrades performance as states from the relevant partition start to be left out of the training set. The benchmark provides an ideal testing ground to investigate these types of behaviours.
4 Discussion and Conclusions
@@
This work provides three main new contributions to the reinforcement learning methodology for solving Markov decision processes and optimal control tasks.
Defining an optimal control problem for a sequence of natural images requires understanding how the problem can be formulated, and what types of solution we can expect. We derived the general conditions for an image to be sufficient for implementing an optimal policy by drawing on an analogy with belief states in partially observed Markov decision processes. The first, and most obvious, condition is that an image must contain all task-dependent information required for the optimal control task at hand so that the cost or reward of each time period can be determined. The second, less obvious, condition is that we must have access to a mechanism for generating new images, where the next image is determined by the current image and the control we choose at the current time period. An image generator such as this is necessary for capturing the dynamics of moving through an environment under a sequential decision process when each state of the environment is an image.
This work also contributed a novel reinforcement learning benchmark for comparing different image representations. The benchmark scales easily so that large numbers of states can be considered for both finite and infinite horizons. This was necessary to convincingly demonstrate the scalability of efficient image representations to large tasks with many states. The benchmark implements a particularly simple image generator, making use of natural image patches and generative stochastic dynamics motivated by target tracking in natural video. This approach allows a simple cost per time period to be implemented, and optimal policies to be easily distinguished from suboptimal policies. More generally, a useful practical image generator might comprise a camera attached to an autonomous vehicle with kinematic controls, or a future generative AI technology that can generate natural images or video frames given some partial information about the environment.
The main contribution of this work was to propose and demonstrate the effectiveness of overcomplete sparse codes as natural image representations that are computationally efficient for optimal control. This work is a significant extension of previous results in this area achieved without a scalable benchmark (Loxley, 2021). When a linear network is used to find optimal policies for image sequences, computationally efficient image representations were shown to lead to faster network training with fewer least-squares iterations, and to minimal-sized networks for storing the input-output associations in training sets. Therefore, these representations use fewer computational resources to learn and evaluate optimal policies. As a result, it now becomes possible to consider larger tasks with more states. The best result presented here was finding an optimal policy for a task with 21675 natural images at each time period, with each image of size 361 pixels. This result was achieved using a ×$\times 64$ sparse code (at 94% efficiency), with the size of the task approximately three orders of magnitude larger (in base 4) than could be achieved using any complete code; for example, one generated using independent component analysis.
Computationally efficient representations were shown to be both overcomplete and sparse. By considering an overcomplete natural image, it was demonstrated that overcompleteness alone is not sufficient for increasing the storage capacity of a linear network. Similarly, sparseness alone is not sufficient. Only by combining both properties to get an overcomplete sparse code is this computational efficiency exhibited. The underlying reason is due to the form of the design matrix for the least-squares problem of network training. Provided the number of states is not larger than the size of the overcomplete sparse code, the design matrix is approximately square. Sparsity then approximately decorrelates neighbouring pixels in the input image, maximizing the rank of the design matrix. In turn, this maximizes the storage capacity of the network and leads to a more favourable condition number for the Hessian matrix.
An interesting question now becomes: what is the limit to overcompleteness for sparse image representations? In this work, we looked at ×$\times 4$, ×$\times 16$, and ×$\times 64$ overcomplete sparse codes. Might it be possible to continue to increase the degree of overcompleteness, and thereby continue to increase the storage capacity of a linear network for an input of fixed size? Achieving this may require a more sophisticated approach to sparsity. Here, we employed a very simple method for generating overcomplete sparse codes. It is possible to achieve sparser codes with alternative algorithms, such as those described in Olshausen (2013), and Rehn and Sommer (2007).
In this work, we also used the benchmark to explore affects due to the length of the horizon, alternative sampling strategies, and the partition of states into greedy optimal and suboptimal policies. We showed it was not necessary to train on all states in order to find suboptimal greedy policies. We also showed that for some initial states, suboptimal greedy policies only develop as the length of the horizon increases. An infinite horizon analysis confirmed the existence of stationary policies that are suboptimal. The benchmark provides an ideal testing ground to investigate these types of behaviours.
The work presented here is highly relevant to anyone interested in developing efficient solution methods for Markov decision processes or optimal control tasks involving natural image sequences or natural video frames. It may also be relevant more generally for optimal control with other types of high-dimensional data that lead naturally to sparse codes. Sparse codes can be generated by devices with low energy requirements (Fair et al., 2019), and have a much lower computational overhead when compared with competing AI methodologies such as deep learning. This makes overcomplete sparse codes worthy of consideration for optimal control applications.
@@
The method used here to generate overcomplete sparse codes of a natural images is taken from Loxley (2017, 2021), and starts with the (real-valued) two-dimensional (2D) Gabor function given by:
@@
where $k({r}^{\prime})=2\pi/\lambda({r}^{\prime})$; and where ${r}=(i,j)$ and ${r}^{\prime}=(i^{\prime},j^{\prime})$ are discrete two-dimensional coordinates. When the 2D Gabor function is adapted to natural image statistics, the three spatial Gabor parameters $\sigma_{x},\sigma_{y},$ and $\lambda$ are found to be strongly correlated and have heavy-tailed distributions (Loxley, 2017). The joint probability density of these parameter values is approximated using the sampling scheme in Table 4, and described by a Gaussian copula with Pareto marginal distributions.
@@
subject $\text{subject to}\ \ \ a\in\mathbb{R}^{m}.$
@@
so that $\det{(\boldsymbol{A})}=1-\alpha p-\alpha^{3}(1-p)$. For the stationary optimal policy, $\boldsymbol{g}$ is given by $[1,0,0]^{\top}$, and it is easily confirmed that $\boldsymbol{AJ}_{\mu}^{*}=[1,0,0]^{\top}$ is solved by
@@
The stationary greedy policy identified in Sec. 3.1 is given by $\mu^{G}((0,0),(0,0))=(1,0)$, $\mu^{G}((0,-1),(1,1))=(0,1)$, and $\mu^{G}((0,-1),(0,1))=(0,1)$. These states have a cost per time period given by $g((0,0),(0,0))=0$, $g((0,-1),(1,1))=1$, and $g((0,-1),(0,1))=1$. Now $\boldsymbol{g}$ is given by $[0,1,1]^{\top}$. It can then be confirmed that $\boldsymbol{AJ}_{\mu}^{G}=[0,1,1]^{\top}$ is solved by
@@
This ratio describes the gap in the expected total cost between the greedy and optimal policies in the infinite horizon limit, and is useful for comparing with the finite horizon results of Sec 3.1. Some useful values include: 2 (when $p=0$), 5 (when $p=0.75$), and $1/(1-p)$ as $p\rightarrow 1$. Alternatively, for $p=1$, the original expression (27) gives $1/(1-\alpha)$ as $\alpha\rightarrow 1$.

### v2

@@ 4.6 Benchmark Performance with Partial Training Sets
Now we consider situations where only some of the states (image patches) of the benchmark are used to train the network. One possibility would be to focus training on initial states that have a suboptimal greedy policy. However, under an optimal policy states outside this set will be visited. A better approach makes use of the analysis in the previous section to partition the benchmark states into two sets. States with at least one non-negative coordinate for $c_{k}-t_{k}$ in the benchmark correspond to a situation where the tracker is not both behind and to the left of the target. These states are independent of the set of states where the tracker is both behind and to the left of the target: in which case the tracker never reaches the target. When states partition in this way, it is possible to focus training on the relevant partition to achieve optimal performance.
@@
Figure 10: Expected total cost of optimal and greedy policies for each initial state with a suboptimal greedy policy when only 78% of the training set is used (see Text for details). Initial states are indexed from 1 to 672. Fitted value iteration (Fitted VI) remains optimal (Inset). The benchmark horizon is 400 time periods, and $p=0.4$.
@@
In Fig 10, the benchmark is used with a $\times 4$ sparse code leading to 1323 states (a small number of states is used here as it is easier to display). Of these, 1031 states have at least one non-negative coordinate (i.e., 78% of the total number of states). These states are used as the training set. Fitted value iteration is then applied to find optimal policies for the 672 initial states with suboptimal greedy policies. Fig 10 shows these policies remain optimal under the proposed training strategy. However, further decreasing the size of the training set rapidly degrades performance as states from the relevant partition start to be left out of the training set. The benchmark provides an ideal testing ground to investigate these types of behaviours.
5 Conclusion and Discussion
@@
This work investigated optimal control over a sequence of natural images and provides three new contributions to reinforcement learning methodology.
The first consideration is what information to include in a state to allow for optimal control over a sequence of images. We found that an image can be a state provided certain conditions are met. The most obvious condition is that an image must contain all task-relevant information leading to a cost or reward at each time period. On the other hand, a less obvious condition is that any model of the dynamics must be capable of generating images from the environment. The reason is that a controllable Markov chain, or a Markov decision process, provides a mechanism for generating the next state from the current state and a chosen control. When each state is an image, the dynamics of moving through an environment generates a sequence of images.
The next consideration is how to encode an image, or find a suitable image representation, to make optimal control computationally efficient. Previous work suggested this might be possible by making use of sparse codes from visual neuroscience (loxley2). To test this idea we developed a new reinforcement learning benchmark that can be used to compare different image representations for computational efficiency during optimal control. The new benchmark is simple to implement, makes it easy to distinguish between optimal and suboptimal policies, and scales to large numbers of states and long horizons. The infinite horizon limit also leads to useful stationary policies. These properties are necessary to convincingly demonstrate the scaling of image representations and computational efficiency to large tasks with many states.
The main contribution of this work was to demonstrate that representing natural images as overcomplete sparse codes is computationally efficient for finding optimal policies in reinforcement learning. This is a significant result. The best result presented here was finding an optimal policy using a $\times 64$ overcomplete sparse code, allowing for the solution of an optimal control task approximately three orders of magnitude larger (in base 4) than could be solved using a complete code such as one generated using ICA or PCA. The underlying reason for the computational efficiency of an overcomplete sparse code can be related to the form of the linear least squares problem for training the final layer of a neural network. Provided the number of images is no larger than the size of an overcomplete sparse code, the design matrix of the least squares problem is approximately square. A key property of a sparse code is that neighbouring pixels in an input image are approximately decorrelated. Decorrelation increases the rank of the design matrix, and favourably conditions the Hessian matrix for neural network training (loxley2). In turn, this maximizes the storage capacity of the linear network and reduces the number of least squares iterations required for training. Since standard decorrelation techniques do not generalize to overcomplete representations, sparsity and overcompleteness seem to be minimum requirements for increasing neural network storage capacity efficiently. An interesting question then becomes: what is the limit to overcompleteness for a sparse code?
This work also shows deep learning is not necessary for efficient optimal control with natural images. A deep neural network may be able to learn specific features for a specific image dataset to help with optimal control. However, a sparse code only depends on the statistics of natural images, and therefore provides a set of general purpose features that can be used on any image dataset. Of course, deep learning can also be used to learn sparse codes (papyan; li). Our model has the advantage that it is simpler to train and uses less memory than a typical deep neural network. The time complexity of training our model depends upon the solution of two linear least squares problems. The first least squares problem has a $d\times m$ design matrix (for a single image with $d$ pixels and a sparse code of size $m$), leading to $O(dmt_{1})$ for an iterative algorithm such as gradient descent; with $t_{1}$ the number of iterations until convergence. For $n$ images this must be repeated $n$ times, leading to $O(dmt_{1}n)$. The second least squares problem has a $m\times n$ design matrix, leading to $O(mnt_{2})$; with $t_{2}$ the number of iterations until convergence. For optimal control over $N$ time periods this must be repeated $N$ times, leading to $O(mnt_{2}N)$. The overall time complexity of training is therefore $O(mn(dt_{1}+Nt_{2}))$. Both least squares problems are well-conditioned (at least until $n\rightarrow m$, see Figure 7) so the iteration numbers $t_{1}$ and $t_{2}$ are not excessive. Our model only requires memory to store $m(d+1)$ weights. Finally, it should be noted that sparse codes can also be generated directly from hardware using purpose built low-power neuromorphic computer chips (chip).
The original motivation for this work was to investigate possible computational advantages of the efficient coding hypothesis from neuroscience. If neural codes do in fact make use of sparseness and overcompleteness, this work has demonstrated some computational advantages to doing so. In particular, the decorrelating effect of a sparse code increases the speed of learning, while an overcomplete sparse code can additionally take advantage of an increase in the size and number of weights of a network to efficiently increase its memory capacity. These findings provide two possible explanations for the appearance of sparse codes in neuroscience.
@@
The method used here to generate overcomplete sparse codes of a natural images is taken from loxley1, loxley2, and starts with the (real-valued) two-dimensional (2D) Gabor function given by:
@@
where $k({r}^{\prime})=2\pi/\lambda({r}^{\prime})$; and where r=(i,j){r}=(i,j) and ${r}^{\prime}=(i^{\prime},j^{\prime})$ are discrete two-dimensional coordinates. When the 2D Gabor function is adapted to natural image statistics, the three spatial Gabor parameters $\sigma_{x},\sigma_{y},$ and $\lambda$ are found to be strongly correlated and have heavy-tailed distributions [loxley1]. The joint probability density of these parameter values is approximated using the sampling scheme in Table 4, and described by a Gaussian copula with Pareto marginal distributions.
@@
$\text{subject to}\ \ \ a\in\mathbb{R}^{m}.$
@@
so that $\det{(\boldsymbol{A})}=1-\alpha p-\alpha^{3}(1-p)$. For the stationary optimal policy, $\boldsymbol{g}$ is given by $1,0,0]^{\top}$, and it is easily confirmed that $\boldsymbol{AJ}_{\mu}^{*}=[1,0,0]^{\top}$ is solved by
@@
The stationary greedy policy identified in Sec. 4.1 is given by $\mu^{G}((0,0),(0,0))=(1,0)$, $\mu^{G}((0,-1),(1,1))=(0,1)$, and $\mu^{G}((0,-1),(0,1))=(0,1)$. These states have a cost per time period given by $g((0,0),(0,0))=0$, g((0,1),(1,1))=1g((0,-1),(1,1))=1, and g((0,1),(0,1))=1g((0,-1),(0,1))=1. Now $\boldsymbol{g}$ is given by $0,1,1]^{\top}$. It can then be confirmed that $\boldsymbol{AJ}_{\mu}^{G}=[0,1,1]^{\top}$ is solved by
@@
This ratio describes the gap in the expected total cost between the greedy and optimal policies in the infinite horizon limit, and is useful for comparing with the finite horizon results of Sec 4.1. Some useful values include: 2 (when $p=0$), 5 (when $p=0.75$), and 1/(1p)1/(1-p) as $p\rightarrow 1$. Alternatively, for $p=1$, the original expression (26) gives $1/(1-\alpha)$ as $\alpha\rightarrow 1$.